1. ssuming that an oscilloscope displays a Vmax of 5.22V and Vmin of 1.33V. Calculate the percentage of modulation. (5 points)
  2. Three (3) AM broadcast stations are spaced at 18 kHz, beginning at 73 kHz. Each station is allowed to transmit modulating up to 6 kHz. Compute for the upper and lower sidebands of each station and plot it in the frequency domain. (15 points)
  3. Given a Vmax of 7.45V and a modulation index of 0.691, calculate for Vmin. (5 points)
  4. A station is given a carrier frequency of 88 kHz. Having a modulating frequency of 12 kHz, compute for the upper and lower sidebands of the station. (5 points)
1
Expert's answer
2021-10-07T09:03:17-0400

Question 1

Percentage modulation M%=

"\frac {V max-V min}{V max+ V min} ×100"%


"%" = "\frac{5.22-1 .33}{5.22+1.33} ×100" %


= 59.4%

Question 2:

For station 1:

flsb =73-6 =67kHz

fusb = 73+6 = 79kHz

For station 2:

flsb =[73+18]-6=85kHz

fusb =[73+18]+6=97kHz


For station 3:

flsb =(73+18+18)-6=103kHz

fusb = (73+18+18)+6=115kHz



Question 3:

Modulation index M= "\frac{V max-V min}{V max+V min}"


0.691="\frac{7.45-V min}{7.45+V min}"

(0.691×7.45)+0.691(V min)=7.45- V min

1.691V min= 2.3021

V min= 1.3614


Question 4:

Upper side band

=fc+fm=88+12=100kHz

Lower side band= fc-fm=88-12=76kHz


"%"